Please use the values in the resources listed below instead of the textbook values. what are the pH and pOH of a solution of 0.0036 M HCI, which ionizes completely? (Assume Kw-1.01x10-14.) pH49 2.4 pOH 4011.6
Answer #1

Since HCl is a strong acid which completely ionized.

Therefore, [HCl] = [H+] = 0.0036M

pH = -log[H+] = -log[0.0036]

pH = 2.44

pH + pOH =14

pOH = 14 - pH = 14 - 2.44

pOH = 11.56